OpenStudy (anonymous):

Determine whether the scenario below involves permutations or combinations. At a family gathering, 32 people are trying to fit into several cars. Your car holds 6 people. How many different groups of 6 could you take in your car?

OpenStudy (anonymous):

is this a combination question? & if so would u do 32 C 6 ?

OpenStudy (anonymous):

do u know it?

OpenStudy (mary.rojas):

http://answers.yahoo.com/question/index?qid=20111129141443AAOMYOP whats 32 C 6?

OpenStudy (anonymous):

that's the same as doing: 32!/ 6! (32-6)!

OpenStudy (anonymous):

@AwesomeB

OpenStudy (mary.rojas):

Check that last link, I think it will help you!

OpenStudy (anonymous):

no not really.

OpenStudy (mary.rojas):

but isnt it the same question?

OpenStudy (anonymous):

w/ different numbers but they don't really explain it

OpenStudy (anonymous):

I wonder if anyone else knows...

OpenStudy (mary.rojas):

@uri help?

OpenStudy (mary.rojas):

@AccessDenied help?

OpenStudy (mary.rojas):

@nincompoop help?

OpenStudy (mary.rojas):

@satellite73 help?

OpenStudy (mary.rojas):

@hero help?

OpenStudy (mary.rojas):

@radar help?

OpenStudy (mary.rojas):

@precal help?

OpenStudy (mary.rojas):

@agent0smith help?

OpenStudy (mary.rojas):

@music101 does this help? http://www.mathsisfun.com/combinatorics/combinations-permutations.html

OpenStudy (anonymous):

I just want to check if my answer is right..

OpenStudy (anonymous):

i got 906192???

OpenStudy (anonymous):

i think thts correct isn't it?

OpenStudy (anonymous):

I believe this is a combination question, because I don't think they care which of the 6 people sit in which seat. They just ask how many groups of 6 you can have

OpenStudy (anonymous):

yea it is

OpenStudy (anonymous):

Ok, so we'll evaluate 32 C 6, which as you said is equal to this: \[\frac{ 32! }{ 6! \times (32-6)! }=\frac{ 32! }{ 6! \times 26! }\]

OpenStudy (anonymous):

yup so im right

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what abt this?

OpenStudy (anonymous):

Ten elementary school students are eligible to be appointed to two positions: plant caretaker and chalkboard assistant. How many unique arrangements of these two positions are possible?

OpenStudy (anonymous):

is it a permutation q.?

OpenStudy (anonymous):

I don't remember the definition of permutation, but the answer will involve a factorial. (Or part of a factorial.)

OpenStudy (anonymous):

@hero do i just do 10 P 2??

OpenStudy (anonymous):

Just looked up permutation and the answer is yes, it is.

OpenStudy (agent0smith):

is this a combination question? & if so would u do 32 C 6 ? correct. Second one is a permutation because order matters, since the two positions are different.

OpenStudy (anonymous):

yea so wouldn't i just do 10P2?

OpenStudy (agent0smith):

Yes.

OpenStudy (anonymous):

because its a permutation question

OpenStudy (anonymous):

ok so i got 90??

OpenStudy (agent0smith):

Or just 10*9. There's 10 students for the first position, and 9 students left for the second.

OpenStudy (anonymous):

ok thank u :)

OpenStudy (anonymous):

when they say "how many UNIQUE arrangements are possible" can u say that its a permutation?? or..

OpenStudy (agent0smith):

Pretty much. Arrangements = permutation. Arrangement means that order/position matters.

OpenStudy (agent0smith):

Groups = combination.

OpenStudy (anonymous):

ok got it

OpenStudy (agent0smith):

Also, ways = permutation. How many ways can ten people in a running race finish first/second/third = permutation question.

OpenStudy (anonymous):

ok but i don't get this question: Four people from a group of 300,000 radio listeners are eligible for a free vacation. How many different groups of vacationers are possible?

OpenStudy (anonymous):

i did 300,000 C4 but its kinda complicated...

OpenStudy (anonymous):

ill be back in a while but i still need help w/ this problem

OpenStudy (agent0smith):

Er yeah you'll get huge numbers with that one...

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=300%2C000+Choose+4

OpenStudy (agent0smith):

lol, my ipod calculator doesn't even do 300,000! which you need to actually calculate that. Just use the wolfram alpha result.

OpenStudy (agent0smith):

or simplify like so \[\Large 300000 C 4 = \frac{ 300000! }{ 4! (300000- 4)!} = \]\[\Large \frac{ 300000! }{ 4! 299996! } = \frac{ 300000 *299999*299998*299997 }{ 4! }\] and then your calculator can do it.

OpenStudy (anonymous):

so then what would be the answer

OpenStudy (agent0smith):

I posted a link to it above.

OpenStudy (anonymous):

... is that even the answer? or is that the wrong way to do it

OpenStudy (agent0smith):

Try calculating the one i gave above and see if you get the same answer.. this one \[\Large \frac{ 300000 *299999*299998*299997 }{ 4! }\]

OpenStudy (anonymous):

I got an answer with an E at the end of it

OpenStudy (agent0smith):

If you got something like 3.37 E20, that means\[ \large 3.37 \times 10 ^{20} \]

OpenStudy (agent0smith):

E at the end of a number on a calculator just means x10^

OpenStudy (anonymous):

so I would put my answer in as 3.3749325*10^20???

OpenStudy (anonymous):

what is: a_______ is a grouping of k objects taken from a set of n elements?

OpenStudy (anonymous):

is that a COMBINATION?

OpenStudy (agent0smith):

Well you can round it to 2 decimal places... 3.37*10^20. There's a keyword in that question that tells you if it's a combination or permutation... remember these keywords from earlier: arrangements, groups, ways. Scroll back up if you forgot which is combination and permutation.

OpenStudy (anonymous):

nvm I actually need help w/ this:

OpenStudy (anonymous):

For your birthday, your parents take you to dinner and offer to let you have any 2 desserts on the menu. The restaurant, famous for its desserts, lists 10 on its menu. How many different combinations of 2 desserts can you order?

OpenStudy (anonymous):

is it 10C2 or how do u do it

OpenStudy (agent0smith):

Well the C basically means choose. So how many desserts do you have to choose from? 10. And how many do you need to choose? 2.

OpenStudy (anonymous):

oh ok then I got 45

OpenStudy (anonymous):

but I don't know how he got 28...

OpenStudy (agent0smith):

nCr means "from n objects, choose r". Eg 10C2 = from 10 desserts, choose 2. 10C2 = 45.

OpenStudy (anonymous):

ok and when a question says " how many unique itineraries are possible for u to choose", its a permutation right?

OpenStudy (agent0smith):

Yep, sounds like it. An itinerary is like a list, so order does matter.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got 151200

OpenStudy (agent0smith):

Well you didn't post the whole question so idk if that's right...

OpenStudy (anonymous):

oh right :p

OpenStudy (anonymous):

well there's 10 cities u want to go to but u only have time for 6

OpenStudy (anonymous):

that's what the question basically says

OpenStudy (anonymous):

so I used permutation and got that answer(151200)

OpenStudy (agent0smith):

Then 10P6 looks right, as does your answer.

OpenStudy (anonymous):

i don't understand D: i missed the "32 people tring to fit into several cars" question, the "10 elementary students appointed to 2 positions" question, the group of 7 musicians competing to earn 3 spots questions, and the one w/ the "E" answer question :( :(

OpenStudy (anonymous):

i knew id probably miss that last one but not the others...

OpenStudy (agent0smith):

I'm not sure how you missed them, your answers here looked good. At a family gathering, 32 people are trying to fit into several cars. Your car holds 6 people. How many different groups of 6 could you take in your car? This should be 32C6, unless you're one of the 32 people and it's meant to be 31C6. The 10 students should be 10P2. the group of 7 musicians competing to earn 3 spots questions ^This is either 7C3 or 7P3, depending on the actual question. And the one with the 'E' was 3.37*10^20.

OpenStudy (anonymous):

i know D: I'll HAVE to talk to my teacher abt it. i missed 4 questions!!! that's horrible!

OpenStudy (agent0smith):

Well not really horrible, considering your answers appeared correct, based on the questions you posted here.

OpenStudy (anonymous):

yea but the second time i took this its more like a mini-quiz and its like a B-minus...

OpenStudy (anonymous):

my 1st one :( i cant repeat that grade though i don't like anything lower than an A

OpenStudy (anonymous):

hey r u good @ conditional probability though

OpenStudy (agent0smith):

Well you can always check with the teacher, maybe the answer key was wrong. I'm pretty sure your answers were right.

OpenStudy (anonymous):

its not an ans key though, u just submit an attachment and its like a computer scored assignment

OpenStudy (agent0smith):

I mean the computer scoring system might be off. I don't see how any of your answers were wrong.

OpenStudy (anonymous):

yea ill ask my teacher abt it. so r u good @ conditional probablility

OpenStudy (agent0smith):

Yeah, probably.